Help me again please!..another mixture problem....Thank you very much *How much decaffeinated coffee(assume to be 0% caffeine)and 40% caffeine coffee should be mixed to produce 20 mugs of 30% caffeine coffee?

Answer by Alan3354(69443)

How much decaffeinated coffee(assume to be 0% caffeine)and 40% caffeine coffee should be mixed to produce 20 mugs of 30% caffeine coffee?
------------
d = decaf
c = regular
----------
d + c = 20 (total mugs)
0*d + 40*c = (0 + 40)*30 (total caffeine)
40c = 1200
c = 30 mugs
d = 10 mugs